OpenStudy (anonymous):

What is the y coordinate of the y intercept of the line: y = -2x + 11

OpenStudy (anonymous):

11

OpenStudy (ivettef365):

y = mx + b m = slope b = y -int

OpenStudy (anonymous):

Alternatively, you can always find the y intercept of any equation by plugging in x=0 then evaluating for y.

OpenStudy (anonymous):

thanx! can you show me how to solve it guys?

jimthompson5910 (jim_thompson5910):

y = -2x + 11 y = -2(0) + 11 y = 0 + 11 y = 11 So the y-intercept is the point (0,11) therefore, the y coordinate of this point is 11

OpenStudy (anonymous):

thanks Jim

jimthompson5910 (jim_thompson5910):

np